\begin{thebibliography}{99}
\bibitem{Kolmogorov}
Kolmogorov, A. N., {\it "Foundation of the Theory of Probability"},  Chelsea: New York,
NY, USA, 1956. ISBN-10:0821826484. \newline ISBN-13:9780821826485.

\bibitem{Holevo}
Holevo, A. S., {\it Probabilistic and Statistical Aspects of Quantum Theory}, 
North-Holland Publishing Company: Amsterdam, The Netherlands, 1982 \newline
ISBN: 9780444863331/0444863338.

\bibitem{Randomness}
Khrennikov, A., {\it Probability and Randomness. Quantum versus Classical}; World Scientific: Singapore, 2016. ISBN-10:1783267968, ISBN-13:978-1783267965.
\newline {\sl Int. J. Quantum Information}, {\bf 14}, p.1640009 (2016)\newline
https://doi.org/10.1142/S0219749916400098

\bibitem{Sch1935}
Schr\"odinger, E., "Discussion of probability relations between separated systems", {\sl Mathematical Proceedings of the Cambridge Philosophical Society}, {\bf 31 (4)}, pp. 555–563 (1935) %Bibcode:1935PCPS...31..555S. 
\newline doi:10.1017/S0305004100013554. S2CID 121278681.

\bibitem{Sch36}
Schr\"odinger, E.,  "Probability relations between separated  systems", {\sl Mathematical Proceedings of the Cambridge Philosophical Society} {\bf 32 (3)}, pp. 446–452 (1936) %Bibcode:1936PCPS...32..446S. 
\newline doi:10.1017/S0305004100019137. S2CID 122822435.

\bibitem{1}
Nielsen, M. A., and Chuang, I. L., {\it "Quantum Computation and Quantum Information, 10th Anniversary ed."}, University Press: Cambridge, UK, 2010; p. 704 \newline ISBN 978-1-107-00217-3.

\bibitem{Bell}
 Bell, J. S., "On the Einstein--Podolsky--Rosen Paradox" {\sl Physics}, {\bf 1 (3)}, pp. 195–200 (1964) \newline doi:10.1103/PhysicsPhysiqueFizika.1.195.

\bibitem{Khrennikov}
Khrennikov, A., "Against Identification of Contextuality with Violation of the Bell Inequalities: Lessons from Theory of Randomness", {\sl J. Russ. Laser Res.}, {\bf 43}, pp. 48–59 (2022). \newline https://doi.org/10.1007/s10946-022-10022-9
\newline arXiv:2112.03910 [quant-ph] (2021)

\bibitem{AndreevMankoDubna}
Andreev, V. A., Man'ko, V. I., Man'ko, O. V., and Sckhukin, E.  V., "Tomography of Spin States, the Entanglement Criterion, and Bell's Inequalities", {\sl Theor. Math. Phys.}, {\bf 146}, N 1, pp. 140–151 (2006) \newline
https://doi.org/10.1007/s11232-006-0014-8

\bibitem{Andreev2007}
Andreev, V. A., "Generalized Bell inequality and a method for its verification", {\sl Theor. Math. Phys.}, {\bf 152}, pp. 1286–1298 (2007) \newline
https://doi.org/10.1007/s11232-007-0113-1

\bibitem{Andreev2009}
Andreev, V. A., "The correlation Bell inequalities", {\sl Theor. Math. Phys.}, {\bf 158}, pp. 196–209 (2009) \newline https://doi.org/10.1007/s11232-009-0016-4

\bibitem{Entropy2022}
Chernega, V. N., Man'ko O. V., and Man'ko V. I.,
"Entangled Qubit States and Linear Entropy in the Probability Representation of Quantum Mechanics", {\sl Entropy}, {\bf 24(4)}, p. 527 (2022) 
\newline https://doi.org/10.3390/e24040527

\bibitem{ManciniPL}
Mancini, S., Man'ko, V. I., and Tombesi, P., "Symplectic Tomography as Classical Approach to Quantum Systems", {\sl Phys. Lett. A}, {\bf 213}, pp. 1-6 (1996)\newline 
https://doi.org/10.1016/0375-9601(96)00107-7

\bibitem{ManciniFP} 
Mancini, S., Man'ko, V.I., and  Tombesi, P., "Classical-like description of quantum dynamics by means of symplectic tomography", {\sl Found. Phys.}, {\bf 27}, pp. 801–824 (1997) \newline https://doi.org/10.1007/BF02550342

\bibitem{TombesiSemOpt} 
Mancini, S., Man'ko, V. I., and  Tombesi, P., "Wigner function and probability distribution for shifted and squeezed quadratures", {\sl J. Opt. B: Quantum Semiclass. Opt.}, {\bf 7}, p. 615 (1995)\newline
DOI 10.1088/1355-5111/7/4/016

\bibitem{OVMVIMLasRes1997}
Man'ko, O. V., and Man'ko, V. I., "Quantum states in probability representation and tomography, {\sl J. Rus. Laser Res.}, {\bf 18},  pp. 407–444 (1997) \newline 
https://doi.org/10.1007/BF02559668

\bibitem{Mancini2} 
Caponigro, M., Mancini, S., and Man'ko, V. I.,  "A probabilistic approach to quantum mechanics based on "tomograms"", {\sl Fortschritte der Physik}, {\bf 54}, p. 602 (2006) \newline
https://doi.org/10.1002/prop.200610304

\bibitem{Dodonov}
Dodonov, V. V., and Man’ko, V. I., "Positive Distribution Description for Spin States", {\sl Phys. Lett. A}, {\bf 229}, pp. 335–339 (1997)
\newline 
https://doi.org/10.1016/S0375-9601(97)00199-0

\bibitem{JETP}
Man’ko, V. I., and Man’ko, O. V., "Spin State Tomography", {\sl J. Exp. Theor. Phys.}, {\bf 85}, pp. 430–434 (1997) \newline 
https://doi.org/10.1134/1.558326

\bibitem{Bregenze}
Man’ko, O. V., "Tomography of Spin States and Classical Formulation of Quantum Mechanics", in: Gruber, B., Ramek, M. (eds), {\it Symmetries in Science X}, 1998, Springer, Boston, MA, 1998, p. 207 \newline 
 doi:10.1007/978-1-4899-1537-5-%_%
 -13

\bibitem{TerraCunha}
Terra-Cunha, M. O., Man’ko, V.I.,  and Scully, M. O.,  "Quasiprobability and Probability Distributions for Spin-1/2 States", {\sl Found. Phys.
Lett.},  {\bf14}, pp. 103–117 (2001) \newline
https://doi.org/10.1023/A:1012373419313

\bibitem{Weigert} 
Amiet, J. P., and Weigert, S., "Coherent states and the reconstruction of pure spin states", {\sl J. Opt. B: Quantum Semiclass. Opt.}, {\bf 1}, pp.  L5 - L8 (1999)
DOI 10.1088/1464-4266/1/5/101

\bibitem{Paini}
D'Ariano, G. M., Maccone, L., and Paini, M., "Spin tomography", {\sl J. Opt. B: Quantum Semiclass. Opt.}, {\bf 5}, p. 77 (2003)\newline
DOI 10.1088/1464-4266/5/1/311

\bibitem{SCH26}
Schr\"odinger, E., "Quantisierung als Eigenwertproblem (Erste Mitteilung)", {\sl Ann. Phys.}, {\bf 384}, pp. 361–376 (1926)\newline
https://doi.org/10.1002/andp.19263840404

\bibitem{Landau}
Landau, L., "Das D\"ampfungsproblem in der Wellenmechanik",  {\sl Z. Phys.}, {\bf 45}, pp. 430–441 (1927)\newline
DOI 10.1007/BF01343064

\bibitem{vonNeumann}
Neumann, J. von, "Wahrscheinlichkeitstheoretischer Aufbau der Quantenmechanik", {\sl Gött. Nach.}, {\bf 1}, pp. 245–272 (1927)
%Neumann, J. von. "Wahrscheinlichkeitstheoretischer Aufbau der Quantenmechanik", Nachrichten von der Gesellschaft der Wissenschaften zu Göttingen, Mathematisch-Physikalische Klasse 1927 (1927): 245-272. 
\newline http://eudml.org/doc/59230

\bibitem{MankoMarmoSudarshan}
Man’ko, V. I., Marmo, G., Sudarshan, E. C. G., and Zaccaria, F.,  
"Interference and entanglement: an intrinsic approach",
{\sl J. Phys. A: Math. Gen.}, {\bf 35}, N 33, p. 7137 (2002)\newline 
DOI 10.1088/0305-4470/35/33/311

\bibitem{SudarshanJRLR} 
Sudarshan, E. C. G., "Search for Purity and Entanglement",  {\sl J. Russ. Laser Res.}, {\bf 24}, pp. 195–203 (2003) \newline https://doi.org/10.1023/A:1024090024283

\bibitem{MankoMarmoSydarshanJRLR}
Man'ko, V. I., Marmo, G., Sudarshan, E. C. G., and Zaccaria,   F.,  "Entanglement Structure of the Adjoint Representation of the Unitary Group and Tomography of Quantum States", {\sl  J.  Russ. Laser Res.}, {\bf 24}, pp. 507–543 (2003) \newline  https://doi.org/10.1023/B:JORR.0000004166.55179.aa

\bibitem{OVMVIMMarmoJPHYS}
Man'ko, O. V., Man'ko, V. I., and Marmo, G., "Alternative commutation relations, star-produst and tomography", {\sl J. Phys. A}, {\bf 35}, 699-719 (2002) \newline
DOI 10.1088/0305-4470/35/3/315

\bibitem{OVMVIMMarmoPHysScr}
Man'ko, O. V., Man'ko, V. I., and Marmo, G., "Star-product of generalized Weyl symbols on SU(2) group, deformations and tomographic probability distributions", {\sl Phys. Scripta}, {\bf 62}, 446-452 (2000)\newline
DOI: 10.1238/Physica.Regular.062a00446

\bibitem{OVMVIMMarmoKrakow}
Man'ko, O. V., Man'ko, V. I.,and  Marmo, G., "Tomographic map within the framework of star-product quantization",   in: {\it Proceedings of Second International Symposium on Quantum Theory and Symmetries, Krakow, Poland, 2001}, (E.~Kapuschik, A.~Morzela, eds), World Scientific, 126-133 (2001)\newline
https://doi.org/10.1142/9789812777850-%_%
0011

\bibitem{OVMVIMMarmoVitale}
Man'ko, O. V., Man'ko, V. I., Marmo, G.,and  Vitale, P., "Star product, duality and double Lie algebras", {\sl  Phys. Lett. A}, {\bf 360}, 522-532 (2007)\newline 	
https://doi.org/10.1016/j.physleta.2006.08.057

\bibitem{myspie} 
Kuznetsov, S. V., Kyusev, A. V., and Man'ko O. V.,  "Tomographic and statistical properties of superposition states for two-mode systems", {\sl Proc. SPIE}, {\bf 5402}, {\it International Workshop on Quantum Optics 2003}, (20 April 2004); 
\newline https://doi.org/10.1117/12.562263 
\newline	arXiv:1903.10784 [quant-ph]

\bibitem{IzvRAN}
Kuznetsov S. V., Kyusev A.V., Man'ko O.V., and Chernega N. V., "Tomographic and statistical properties of superposition states for two--mode systems", {\sl Bulletin of the Russian Academy of Science, Physics}, {\bf 58}, N 9, p.p. 1239-1243 (2004), Allerton Press\newline
https://www.elibrary.ru/item.asp?id=17641248 (in Russian)

\bibitem{Kuznetsov} 
Kuznetsov, S. V., Man'ko, O. V.,  and  Tcherniega, N. V., "Photon distribution function, tomograms and entanglement in Stimulated Raman Scattering", {\sl J. Opt. B: Quantum Semiclass. Opt}, {\bf 5}, p. 5503 (2003)\newline 
DOI 10.1088/1464-4266/5/4/357

\bibitem{SPIE54}
Kuznetsov, S. V., and Man'ko, O. V. "Entanglement and photon number probability distribution function for Stimulated Raman Scattering", {\sl Proc. SPIE}, {\bf 5402}, pp. 302-313 (2004)\newline 
doi: 10.1117/12.562224
\newline Kuznetsov, S. V ., Kyusev, A. V., and Man'ko, O.V.,  
%Proc. SPIE 5402, p. 314-327 (2004)	
arXiv:1903.10784 [quant-ph]

\bibitem{VRMB} 
Man'ko, O.V., and Tcherniega, N. V., "Tomographic description of Stimulated Brillouin Scattering", {\sl J. Russ. Laser Res.}, {\bf 22}, N 3, pp. 201-218 (2001)\newline 
https://doi.org/10.1023/A:1011304404336

\bibitem{VRMB1} 
 Kuznetsov, S. V., Kusev, A. V., Man'ko, O. V., and Tcherniega, N. V., "Entanglement in the process of  Stimulated Brillouin Scattering", {\sl Acta Physica Hungarica B: Series Quantum Electronics}, {\bf 20/1}, pp. 11-14 (2004)
 https://doi.org/10.1556/APH.20.2004.1-2.4
 
 \bibitem{1998}
 Man'ko, O. V., "Photon distribution function for stimulated Raman scattering",  {\sl Proc. SPIE}, {\bf 4089}, {\it Raman Scattering, (Moscow, 18 February 2000)} (2000)\newline
 https://doi.org/10.1117/12.378116
 
\bibitem{6256}
Man'ko, O. V., and Tcherniega N. V., "Photon-number tomography and symplectic approach to stimulated Raman scattering". {\sl Proc. SPIE}, {\bf 6256W}, {\it ICONO 2005: Ultrafast Phenomena and Physics of Superintense Laser Fields; Quantum and Atom Optics; Engineering of Quantum Information}, pp. 62560W-1--62560W-12 (2006) \newline https://doi.org/10.1117/12.682581

\bibitem{Pathak} 
Giri, S. K., Sen, B., Pathak, A., and Jana, P. C., "Higher-order two-mode and multimode entanglement in Raman processes", {\sl Phys. Rev. A}, {\bf 93}, p. 012340 (2016) 
\newline https://doi.org/10.1103/PhysRevA.93.012340

\bibitem{Perina} 
Pathak, A., K\~repelka, J., and Pe\~rina, J., "Nonclassicality in Raman scattering: Quantum entanglement, squeezing of vacuum fluctuations, sub-shot noise and joint photon–phonon number and integrated-intensity distributions", {\sl Phys.  Lett. A}, {\bf 377}, N 38, pp. 2692-2701 (2013)
\newline
https://doi.org/10.1016/j.physleta.2013.07.046
 
\bibitem{Rohith} 
Rohith, M., and Sudheesh, C., "Visualizing revivals and fractional revivals in a Kerr medium using an optical tomogram", {\sl Phys. Rev. A}, {\bf 92}, p. 053828 (2015)\newline 
https://doi.org/10.1103/PhysRevA.92.053828

\bibitem{Rohith1} 
Rohith, M., and Sudheesh, C., "Signatures of entanglement in an optical tomogram," {\sl JOSA B}, {\bf  33}, pp. 126-133 (2016)\newline 
https://doi.org/10.1364/JOSAB.33.000126 

 \bibitem{mySudarshan}
% \bibitem{separcrit}
 Man'ko, O. V.,Man'ko, V. I., Marmo, G., Shaji, A., Sudarshan, E. C. G., and Zaccaria, F., "Partial positive scaling transform and a separability criterion", {\sl Phys.~Lett. A}, {\bf 339},  pp. 194-206 (2005)
 \newline
 https://doi.org/10.1016/j.physleta.2005.02.070
  
 \bibitem{heatingmap}
 Man'ko, O. V., "Heating map in classical and quantum mechanics", {\sl J. Russ. Laser Res.}, {\bf 26}, pp. 109–115 (2005)\newline
 https://doi.org/10.1007/s10946-005-0010-8
 
\bibitem{Facchi} 
Facchi, P., Ligab\'o, M., and Solimini, S., "Tomography: mathematical aspects and applications," {\sl  Phys. Scr.}, {\bf 90}, N 7, p. 074007 (2015)\newline
DOI 10.1088/0031-8949/90/7/074007

\bibitem{Polkovnikov}
Claeyes., P. W., and Polkovnikov, A.,  
"Quantum eigenstates from classical Gibbs distributions"
{\sl SciPost Phys.}, {\bf 10}, p. 014 (42pp.) (2021)
\newline doi: 10.21468/SciPostPhys.10.1.014

\bibitem{Khrennikov5} 
Khrennikov, A., {\sl Physica A: Statistical Mechanics and its Applications}, {\bf 393}, 207 (2014)

\bibitem{SocialLaser}
Khrennikov., A., {\it Social Laser}, Jenny Stanford Publ. 280 pp.(2020) \newline 
ISBN 978-981-4800-83-9

\bibitem{KhrennikovAlo} 
Khrennikov, A., and Alodjants, A., "Classical (Local and Contextual) Probability Model for Bohm–Bell Type Experiments: No-Signaling as Independence of Random Variables", {\sl Entropy}, {\bf 21(2)}, p. 157 (2019) \newline 
https://doi.org/10.3390/e21020157

\bibitem{Khrennikov1a}
Khrennikov., A., "Quantum postulate vs. quantum nonlocality: on the role of the Planck constant in Bell’s argument", {\sl Found. Phys.}, {\bf 51}, p. 16 (12pp) (2021)
\newline  https://doi.org/10.1007/s10701-021-00430-3

\bibitem{Marmo}
Ciaglia, F. M., Di Cosmo, F., Ibort, A., and Marmo, G., "Evolution of Classical and Quantum States in the Groupoid Picture of Quantum Mechanics", 
{\sl Entropy}, {\bf 22(11)}, p. 1292 (2020)
\newline https://doi.org/10.3390/e22111292

\bibitem{Prud}
Przhiyalkovskiy, Y. V., "Continuous Measurements in Probability Representation of Quantum Mechanics", {\sl 
Proc. Steklov Inst. Math.}, {\bf 313}, pp.193–202 (2021) \newline https://doi.org/10.1134/S0081543821020188
\newline arXiv:2101.07568v1 [quant-ph] 

\bibitem{FacchiLigabo63}
Facchi, P., and Ligab\'o, M., "Classical and quantum aspects of tomography", in:  M. Asorey, J. Clemente-Gallardo, E. Martínez, J. F. Cariñena (Ed.) {\it Proceedings of XVIII International Fall Workshop on Geometry and Physics, Benasque, Spain, 2009}, {\it AIP Conference Proceedings}, {\bf 1260} p. 3 (2010)
\newline https://doi.org/10.1063/1.3479322

\bibitem{Elze} 
Elze, H.-T., Gambarotta, G., and Vallone, F., "General Linear Dynamics—Quantum, Classical or Hybrid", {\sl J. Phys.
Conf. Ser.}, {\bf 306}, p. 012010. (2011)
\newline DOI 10.1088/1742-6596/306/1/012010

\bibitem{Stornaiolo62}
Stornaiolo, C., "Emergent classical universes from initial quantum states in a tomographical description",  
{\sl Int. J. Geom. Meth. in Modern Phys.}, {\bf.17(11)}, p. 2050167 (2020)\newline
DOI: 10.1142/S0219887820501674

\bibitem{Stornaiolo62a} 
Stornaiolo, C., "Tomographic Represention of Quantum and Classical Cosmology", in: C. Moreno González, J. Madriz Aguilar, L. Reyes Barrera (Ed.), {\it Accelerated Cosmic Expansion. Proceedings of the Fourth International Meeting on Gravitation and Cosmology}, {\sl Astrophysics and Space Science Proceedings}, {\bf 38} pp. 1155-1157 (2014) Springer
 \newline 
doi: 10.1142/9789814623995-%_%
0107

\bibitem{M} 
Berra–Montiel, J.;  and Molgado, A. "Tomography in loop quantum cosmology", {\sl Eur. Phys. J. Plus}, {\bf 137}, p. 283 (2022)\newline 
DOI: 10.1140/epjp/s13360-022-02504-1

\bibitem{Gouson}
de Gosson, M. A., 
"Quantum Harmonic Analysis of the Density Matrix", {\sl Quanta},  {\bf 7}, p. 74 (2018)
\newline https://doi.org/10.12743/quanta.v7i1.74

\bibitem{Gouson1}
Gosson, M. A., and de Gosson, C.,
"Symplectic Polar Duality, Quantum Blobs, and Generalized Gaussians" 
{\sl Symmetry}, {\bf 14(9)}, p. 1890 (2022) \newline https://doi.org/10.3390/sym14091890

\bibitem{Gouson2}
Gosson, M. A., "Symplectic Radon Transform and the Metaplectic Representation", {\sl Entropy}, {\bf 24(6)}, p. 761 (2022) \newline 
https://doi.org/10.3390/e24060761

\bibitem{Foukzon} 
Foukzon, J., Potapov, A.  A., Menkova, E., and Podosenov, S. A. "A New Quantum-Mechanical Formalism Based on the Probability Representation of Quantum States," (2016) viXra:1612.0298. 

\bibitem{Belolipechkii}
Chernega, V. N., Belolipetskiy, S. N.,  Man'ko, O. V., and  Man'ko, V. I., "Probability representation of quantum mechanics and star product quantization", {\sl J. of Physics: Conf. Ser.}, {\bf 1348}, {\it Irreversible Processes in Nature and Technics 29–31 January 2019, Moscow, Russian Federation}, p. 012101 (2019) \newline 
DOI 10.1088/1742-6596/1348/1/012101

\bibitem{Bazrafkan} 
Bazrafkan, M. R., and Nahvifard, E., "Stationary perturbation theory in the probability representation of quantum mechanics", {\sl J. Russ. Laser Res.}, {\bf 30}, pp. 392–403 (2009) \newline
https://doi.org/10.1007/s10946-009-9079-9

\bibitem{Filinov}
Filinov, V. S., Schubert, G., Levashov,P., Bonitz, M., Fehske, H., Fortov, V. E., and Filinov,A. V., "Center-of-mass tomographic approach to quantum dynamics", {\sl Phys. Lett. A}, {\bf 372},  p. 5064 (2008) \newline 
https://doi.org/10.1016/j.physleta.2008.05.053

\bibitem{Chernega1}
Chernega, V. N., Man'ko, O. V., and Man'ko, V. I.,
"Triangle Geometry of the Qubit State in the Probability Representation Expressed in Terms of the Triada of Malevich’s Squares", {\sl J. Rus. Laser Res.}, {\bf 38}, N 2, pp. 141–149 (2017) \newline https://doi.org/10.1007/s10946-017-9628-6

\bibitem{Chernega2}
Chernega, V. N., Man'ko, O. V., and Man'ko, V. I., "Probability Representation of Quantum Observables and Quantum States", 
{\sl J. Rus. Laser Res.}, {\bf 38}, N 4, pp. 324-333 
(2017)\newline 
https://doi.org/10.1007/s10946-017-9648-2

\bibitem{Chernega3} 
Chernega, V. N., Man'ko, O. V., and Man'ko, V. I., 
"Triangle Geometry for Qutrit States in the Probability Representation", {\sl J. Rus. Laser Res.}, {\bf 38}, N 5, pp. 416–425 %141 
(2017)\newline https://doi.org/10.1007/s10946-017-9662-4

\bibitem{Chernega4}
Chernega, V. N., Man'ko, O. V., and Man'ko, V. I.,
"God Plays Coins or Superposition Principle for Classical Probabilities in Quantum Suprematism Representation of Qubit States", {\sl J. Rus. Laser Res.}, {\bf 39}, N 2, pp. 128–139 
(2017)\newline https://doi.org/10.1007/s10946-018-9699-z

\bibitem{Chernega5}
Chernega, V. N., Man'ko, O. V., and Man'ko, V. I., 
"Quantum suprematism picture of Triada of Malevich's squares for spin states and the parametric oscillator evolution in the probability representation of quantum mechanics",
{\sl J. Phys.: Conf. Ser. }, {\bf1071}, p. 012008 (2018)
\newline DOI 10.1088/1742-6596/1071/1/012008

\bibitem{Chernega6}
Chernega, V. N., Man'ko, O. V., and Man'ko, V. I., "Correlations in a system of classical-like coins simulating spin-1/2 states in the probability representation of quantum mechanics",{\sl  Eur. Phys. J. D}, {\bf 73}, p. 10 (2019). https://doi.org/10.1140/epjd/e2018-90487-9

\bibitem{Chernega7}
Chernega, V. N., Man'ko, O. V., and Man'ko, V. I., 
"Probability Representation of Quantum States as a Renaissance of Hidden Variables— God Plays Coins", 
{\sl J. Rus. Laser Res.}, {\bf 40}, N 2, pp. 107–120  
(2019) \newline 
https://doi.org/10.1007/s10946-019-09778-4

\bibitem{Chernega8}
Chernega, V. N., Man'ko, O. V., and Man'ko, V. I., 
"Probability Representation of Photon States and Tomography",
{\sl J. Rus. Laser Res.}, {\bf 40}, N 6, pp. 503–514   
(2019) \newline https://doi.org/10.1007/s10946-019-09831-2

\bibitem{Miroshnichenko}
Miroshnichenko, G. P., "CQED Quantum Tomography of a Microwave Range", arXiv:1510.03155 [quant-ph] (2015)\newline 
https://doi.org/10.48550/arXiv.1510.03155

\bibitem{Koczor}
Koczor, B., Zeier, R., and Glaser, S. J., "Continuous Phase-Space Representations for Finite-Dimensional Quantum States and their Tomography", {\sl Phys. Rev. A}, {\bf 101}, p. 022318 (2020) \newline 
https://doi.org/10.1103/PhysRevA.101.022318
\newline 
arXiv:1711.07994v3 [quant-ph] (2018)

\bibitem{23}
Toninelli. E., Ndagano, B., Valles, A., and Forbes, A.,  "Concepts in quantum state tomography and classical implementation with intense light: a tutorial" {\sl  Advances in Optics and Photonics}, {\bf 11(1)}, p. 67--134 (2019) \newline
https://doi.org/10.1364/AOP.11.000067

\bibitem{12}
Almarashi, A. M., Abd-Elmougod, G. A., and Raqab, M. Z.,  "Quantum Extropy and Statistical Properties of the Radiation Field for Photonic Binomial and Even Binomial Distributions", 
{\sl J. Russ. Laser Res.}, {\bf 41}, pp. 334-343 (2020)\newline 
https://doi.org/10.1007/s10946-020-09883-9

\bibitem{Leon}
Leon, R. C. C., Yang, C. H., Hwang, J. C. C., Lemyre, J. C.,  Tanttu, T., Huang, W., Huang, J. H., Hudson, F. E., Itoh, 
K. M., Laucht, A., Pioro-Ladri\'ere, M., Saraiva, A., and Dzurak,  A. S., "Bell-state tomography in a silicon many-electron artificial molecule", {\sl Nat. Commun.},{\bf 12}, p. 3228 (2021) \newline https://doi.org/10.1038/s41467-021-23437-w
\newline 
arXiv:2008.03968v1 (2020)
\newline 
https://doi.org/10.48550/arXiv.2008.03968

\bibitem{Kiktenko1} 
Kiktenko, E. O., Fedorov, A. K., Man'ko, O. V. and Man'ko, V. I., "Multilevel superconducting circuits as two-qubit systems: Operations, state preparation, and entropic inequalities", {\sl  Phys. Rev. A}, {\bf 91}, p. 042312 (2015)\newline 
DOI:https://doi.org/10.1103/PhysRevA.91.042312

\bibitem{myJos} 
 Man'ko, O.V., "Quantum tomography of current and voltage states in nanoelectriccircuits", {\sl AIP Conference Proc.}, {\bf 1424}, p. 221-233 (2012) \newline
 https://doi.org/10.1063/1.3688974

\bibitem{Mynonlinear} 
  Man'ko, O. V., "Symplectic tomography of nonlinear coherent states of a trapped ion", {\sl Phys. Lett. A}, {\bf 228}, pp. 29-35 (1997) \newline 
 https://doi.org/10.1016/S0375-9601(97)00091-1
  
 \bibitem{cat}
 Man'ko, O. V. , "Symplectic tomography of Sch\"odinger cat states of a trapped ion", in {\it Proc. of "New developments pon Fundamental Problems in quantum physics, Oviedo, Spain, 1996, Foundamental Theories of Physics}, Kluwer Publ. M. Ferrero, A. van der Merwe Eds.p.225-229 (1997) 
DOI: 10.1007/978-94-011-5886-2-%_%
29 
 
  \bibitem{3736}
 Man'ko, O. V., "Optical tomography and measuring quantum states of an ion in a Paul and in a Penning trap",
{\sl Proc. SPIE}, {\bf 3736}, {\it ICONO '98: Quantum Optics, Interference Phenomena in Atomic Systems, and High-Precision Measurements, (25 February 1999)} (1999)\newline  https://doi.org/10.1117/12.340153
 
 \bibitem{KronigPenneydamped} 
 Chernega, V. N., and Man'ko, O. V., "A Damped Oscillator with a $\delta$-Kicked Frequency in the Probability Representation of Quantum Mechanics", {\sl J. Russ. Laser Res.}, {\bf 39}, N 4, pp. 411-419 (2018) \newline 
 https://doi.org/10.1007/s10946-018-9735-z
 
 \bibitem{ActaPhysHung} 
 Klimov, A. B., Man'ko, O. V., Man'ko, V. I., Smirnov, Yu. F., and Tolstoy, V. N., "Tomographic representation of spin and quark states", 
 {\sl J. Phys. A: Math. Gen.}, {\bf 35}, p. 6101 (2002) \newline
 DOI 10.1088/0305-4470/35/29/312
 
 \bibitem{2004JRLR} 
 Man'ko, O. V., and Man'ko, V. I., "Probability-Representation Entropy for Spin-State Tomogram", 
  {\sl J. Russ. Laser Res.}, {\bf 25}, pp. 115-122 (2004)\newline 
 https://doi.org/10.1023/B:JORR.0000018981.06864.12
 
 \bibitem{JRLR2001} 
 Man'ko, O. V., and Man'ko, V. I., "Wave Function in Classical Statistical Mechanics", {\sl J. Russ. Laser Res.}, {\bf 22}, pp. 149-167 (2001)\newline https://doi.org/10.1023/A:1011360006073
 
 \bibitem{hlimit} 
 Man'ko, O. V., and Man'ko, V. I., "Classical Mechanics Is not the $\hbar\rightarrow 0$ Limit of Quantum Mechanics", {\sl J. Russ. Laser Res.}, {\bf 25}, p. 477-492 (2004) \newline
 https://doi.org/10.1023/B:JORR.0000043735.34372.8f
 
 \bibitem{Pilyavets} 
 Man'ko, O. V., Man'ko, V. I., Pilyavets, O.V., "Probability representation of classical states", {\sl J. Rus. Laser Res.},  v. 26, N 6, p. 429-444 (2005) \newline 
https://doi.org/10.1007/s10946-005-0045-x

 \bibitem{Ibort} 
 Asorey, M., Ibort, A., Marmo G., and Ventriglia, F., "Quantum Tomography twenty years later", {\sl  Phys. Scr.}, {\bf 90}, p. 074031 (2015) \newline 
 DOI: 10.1088/0031-8949/90/7/074031
 
 \bibitem{Marmo3} 
 Man'ko, V. I., Marmo, G., Simoni, A., Stern, A., and Ventriglia, F., "An Introduction to the tomographic picture of quantum mechanics", {\sl Phys. Lett. A}, {\bf 343}, p.  251 (2005)  \newline
 DOI: 10.1088/0031-8949/79/06/065013

\bibitem{Schleih} 
Ullinger, F., Zimmermann, M., and Schleich, W. P., "The logarithmic phase singularity in the inverted harmonic oscillator",  {\sl  AVS Quantum Sci.}, {\bf 4(2)}, p. 024402 (2022) \newline 
DOI: 10.1116/5.0074429

\bibitem{12a}		
Subramanyan, V., Hegde, S. S., Vishveshwara, S., and Bradlyn, B., "Physics of the Inverted Harmonic Oscillator: From the lowest Landau level to event horizons", {\sl Annals Phys.}, {\bf 435}, p. 168470 (2021)  %e-Print: 2012.09875 • 
DOI: 10.1016/j.aop.2021.168470	

\bibitem{myEntropy2023}
Man'ko, V. I., and Man'ko, O.V., "Inverted Oscillator Quantum States in the Probability Representation", {\sl Entropy}, {\bf 25(2) }, p.217 (2023) \newline 
https://doi.org/10.3390/e25020217

\bibitem{TeretenkovNosal}
Nosal’, I. A., and Teretenkov, A. E., "Exact Dynamics of Moments and Correlation Functions for GKSL Fermionic Equations of Poisson Type", {\sl Math. Notes}, {\bf 108}, p. 911-925 (2020)\newline 
https://doi.org/10.1134/S0001434620110358

\bibitem{Teretenkov}
Teretenkov, A. E., "Irreversible quantum evolution with quadratic generator: Review", {\sl Infin. Dimens. Anal. Quantum Probab. Relat. Top.}, {\bf 22}, N 4, p. 1930001, 27 pp. (2019) 
\newline https://doi.org/10.1142/S0219025719300019

\bibitem{ManciniPaulieq}
Mancini, S., Man'ko, O. V., Man'ko, V. I., and Tombesi, P., "The Pauli equation for probability distributions", {\sl J. Phys. A: Math. Gen.}, {\bf 34}, p. 3461. (2001)
\newline DOI 10.1088/0305-4470/34/16/314
%\bibitem{Korennoy} 
%Korennoy Ya A 2018 Gauge-independent Husimi Functions of Charged Quantum Particles in the Electro-magnetic Field. 
%arXiv:1806.06443 [quant-ph]
 
\bibitem{Entropy2021}
Man'ko, O. V., and Man'ko, V. I., 
"Probability Representation of Quantum States", {\sl Entropy}, 
{\bf 23(5)}, p. 549 (2021) \newline
https://doi.org/10.3390/e23050549 

\bibitem{Kishore} 
Thapliyal, K., Banerjee, S., and Pathak, A., "Tomograms for open quantum systems: In(finite) dimensional optical and spin systems",  {\sl Annals of Physics}, {\bf 366}, pp. 148-167 (2016)
\newline https://doi.org/10.1016/j.aop.2016.01.010

\bibitem{Arhipov} 
Arkhipov, A. S., and Lozovik, Yu. E., "New method of quantum dynamics simulation based on the quantum tomography", {\sl Phys. Lett. A}, {\bf 319}, pp.  217-224 (2003)\newline 
doi:10.1016/j.physleta.2003.09.084

\bibitem{center}
Arkhipov, A. S.,  Lozovik, Yu. E., and Man'ko, V. I., "Center of mass tomography for reconstructing quantum states of multipartite systems",  {\sl Phys. Lett.A}, {\bf 328}, N 6, pp. 419-431 (2004) \newline  
doi: 10.1016/j.physleta.2004.06.039

\bibitem{center1}
Man’ko, O.V., and  Man’ko, V.I., 
"Fidelity in the center-of-mass tomography", {\sl J. Russ. Laser Res.}, {\bf 31}, pp. 608–613 (2010)\newline https://doi.org/10.1007/s10946-010-9181-z

\bibitem{center2}
Dudinets, I.V., and  Man’ko, V.I., "Center-of-Mass Tomography and Wigner Function for Multimode Photon States", {\sl Int. J. Theor. Phys.}, {\bf 57}, pp. 1631–1644 (2018) \newline  https://doi.org/10.1007/s10773-018-3690-x

\bibitem{Jordan}
Jordan, P., "Zusammenhang der symmetrischen und linearen Gruppen und das Mehrkorperproblem", 
{\sl Z. Phys.}, {\bf 94}, pp. 531-535 (1935)\newline 
doi:  10.1007/BF01330618

\bibitem{schwinger} Schwinger, J.: "On angular momentum", in {\it Quantum Theory of Angular Momentum, eds. L.C. Biedenharn and H. Van Dam}, Academic, 1965.

\bibitem{yesterday1}
Man'ko, M. A., and Man'ko, V. I. , "Hermite Polynomial Representation of Qubit
States in Quantum Suprematism Picture",
In: Marmo, G., Martín de Diego, D., Muñoz Lecanda, M. (eds) {\it Classical and Quantum Physics}, Springer Proceedings in Physics, vol 229. Springer, Cham. (2019) \newline  https://doi.org/10.1007/978-3-030-24748-5-%_%
16

\bibitem{yesterday2}
Lemeshevskiy, D. B., Man’ko, V. I. "Hermite polynomial representation of the spin states", {\sl J.
Russ. Laser Res.}. {\bf 34}, N 2, pp. 175-184 (2013)\newline 
https://doi.org/10.1007/s10946-013-9340-0

\bibitem{yesterday3}
Man'ko, M. A., "Unitary transforms of probability distributions describing quantum states of two-level systems",
{\sl J. Russ. Laser Res.}, {\bf 43}, pp. 645–652 (2022)\newline  https://doi.org/10.1007/s10946-022-10091-w

\end{thebibliography}

\documentstyle[12pt]{article}

\begin{document}
\begin{center}
{\Large{ \bf Entangled probability distributions}}

 {\bf
Vladimir N. Chernega,$^1$ Olga V. Man'ko,$^{2,3}$ Vladimir I.
Man'ko$^{2,4}$}\\
\medskip
{\it
$^1$Russian University of Transport (MIIT), Moscow, Russia\\
\smallskip
$^2$Lebedev Physical Institute, Russian Academy of Sciences\\
Leninskii Prospect 53, Moscow 119991, Russia\\
\smallskip
$^3$Bauman Moscow State Technical University\\
The 2nd Baumanskaya Street 5, Moscow 105005, Russia\\
\smallskip
$^4$Moscow Institute of Physics and Technology (State University)\\
Institutskii per. 9, Dolgoprudny, Moscow Region 141700, Russia}\\
\end{center}
\smallskip
$^*$Corresponding author e-mail:~~~mankoov\,@lebedev.ru

\begin{abstract}\noindent
Concept of entangled probability distribution  of several random variables  is introduced.  These probability distributions describe multimode quantum states in probability representation of quantum mechanics. Example of entangled probability distribution is considered.
\end{abstract}
\medskip

\noindent{\bf Keywords:}
probability distribution, entanglement, tomograms.

\section{ Introduction.}
The foundations of conventional probability theory were considered long ago and are described , e.g. in books  \cite{Kolmogorov,Holevo,Randomness}. The probability theory is necessary to be used to describe random process \cite{Randomness} both classical and quantum. Also to  construct the theory of quantum phenomena like quantum mechanics, quantum statistics, quantum field theory one needs to apply all the aspects of the conventional probability theory.The aim of our work is to study the possibility to use the problem related to back action of the results obtained in quantum mechanics to clarify some aspects which are new in conventional probability theory. The point is that in quantum mechanics new interesting phenomenon called entanglement of quantum system states was found \cite{Sch1935,Sch36}. It is special types of correlations between the subsystems in multiparticle systems. For example, the review of the  entanglement phenomenon in information theory was discussed in \cite{1}. 
The phenomenon of entanglement is related to the properties of quantum systems like Bell inequalities \cite{Bell}. For example, quantum properties reflected from Bell inequalities were discussed in \cite{Khrennikov,AndreevMankoDubna,Andreev2007,Andreev2009,Entropy2022}.  On the other hand recently it was clarified that all quantum states of physical systems can be described by conventional probability distributions using the introduced probability representation of quantum mechanics  \cite{ManciniPL,ManciniFP,TombesiSemOpt, OVMVIMLasRes1997, Mancini2}. The spin tomography was introduced in \cite{Dodonov,JETP} and discussed in \cite{Bregenze,TerraCunha,Weigert,Paini}. 
In \cite{AndreevMankoDubna} in the frame of probability representation of  quantum mechanics the spin tomography was applied to Bell's inequality problem, and was shown that the  satisfaction or violation of Bell's inequalities can be understood as properties of tomographic functions for joint probability distributions for two spins. 

In the probability representation entanglement phenomenon is also associated with the specific properties of the quantum states which may be considered as properties determined by the specific properties of used probability distributions which we call entangled probability distributions. In our work we introduce autonomous notion of the entangled probability distribution and consider their properties on few examples induced by the existed entanglement phenomena in quantum mechanics like these phenomena in oscillator system states. 

In quantum mechanics very important phenomenon is the superposition principle. The superposition principle of wave functions \cite{SCH26} can be discussed using density matrices describing quantum states introduced by \cite{Landau,vonNeumann}. The nonlinear addition of density matrices describing quantum states corresponds to the linear combination of the wave function  \cite{MankoMarmoSudarshan,SudarshanJRLR,MankoMarmoSydarshanJRLR}.  We will use the map of tomograms and density matrices \cite{OVMVIMMarmoJPHYS,OVMVIMMarmoPHysScr,OVMVIMMarmoKrakow,OVMVIMMarmoVitale} to study properties  corresponding to entanglement phenomenon. The tomograms and the entanglement phenomenon in the two mode squeezed states and two-mode even and odd coherent states was considered in \cite{myspie,IzvRAN}.  The application of the symplectic tomography scheme to the  Stimulated Raman Scattering (SRS) and to the Stimulated Brillouin Scattering was done in \cite{Kuznetsov,SPIE54,VRMB,VRMB1} and the entanglement in these processes were discussed.  The classical propagator of SRS was introduced in \cite{1998}. The tomographic entropy, information, entanglement phenomenon and quantum distances between quantum states of photon--phonon mode of SRS in the probability representation were discussed in \cite{6256}. 

The quantum entanglement in the Raman Scattering was also investigated in \cite{Pathak,Perina}. It is necessary to add that optical tomograms of the time-evolved states generated by the evolution of different kinds of initial wave packets in a Kerr medium and optical tomograms of maximally entangled states generated at the output modes of a beam splitter was theoretically studied in \cite{Rohith}, \cite{Rohith1}. The entanglement criterion in probability representation was suggested in \cite{mySudarshan,heatingmap}.
In \cite{Facchi} a review of the Radon transform and the instability of the tomographic reconstruction process were discussed. 
In \cite{Polkovnikov} it was shown that in classical mechanics can be introduced  the Hermitian operators and the concepts of classical mechanics can be formulated using the inverse Wigner--Weyl transform of the classical probability distributions. In \cite{Khrennikov5} the attempts to study the problems of complex biosystems composed of many subsystems and in \cite{SocialLaser} the attempts to model social process using the language of quantum mechanics, thermodynamics, quantum information and field theory were done. In \cite{KhrennikovAlo} a review of classical probability representations of quantum states and observables is done and in \cite{Khrennikov1a} quantum nonlocality is discussed. New fundamental aspects of quantum mechanics based on groupoid approach are investigated in \cite{Marmo}. Development of the tomographic probability distribution method and different aspects of its applications are presented in \cite{Prud}. The classical and quantum aspects of tomography were discussed in \cite{FacchiLigabo63} and classical--quantum dynamics were considered in \cite{Elze}. The probability representation of quantum states is used recently to study properties of cosmology \cite{Stornaiolo62,Stornaiolo62a,M}. The review of the harmonic analyses of the density matrix properties including the method of symplectic tomography is given in \cite{Gouson,Gouson1,Gouson2}. The methods of star-product, tomography and probability representation of quantum mechanics were applied to different problems of quantum phenomena in   \cite{Foukzon,Belolipechkii,Bazrafkan,Filinov}. The geometrical interpretation of spin state in probability representation with the help of Malevich's squares is done in  \cite{Chernega1,Chernega2,Chernega3,Chernega4,Chernega5,Chernega6,Chernega7}.  

We give a review of the tomographic probability representation of quantum mechanics. In \cite{Chernega8} the formalism of quantum states and quantum observables using the formalism of standard probability distributions and classical-like random variables is presented and the evolution equation and energy spectra in the form of equations for probability distributions determining  the coherent and number states of photons are obtained.

Applications of quantum tomographic approach to different kinds of experiments as well as theoretical researches were discussed in \cite{Miroshnichenko,Koczor,23,12,Leon}.
The tomographic causal analysis of two-qubit states was done and theoretically was studied operations with a four-level superconducting circuit as a two-qubit system in \cite{Kiktenko1}. The tomograms of the state of a quantum circuit and of the Josephson junction were considered in \cite{myJos}.  The probability representation of the nonlinear coherent states of an ion in the Paul trap was introduced in \cite{Mynonlinear}, the Schr\"odinger cat states of a trapped ion in the Paul trap were discussed in \cite{cat}, the tomograms of an ion in the Paul and Penning traps were considered in \cite{3736}.  The tomograms of the states of the damped oscillator with Kronig--Penney excitation were considered in \cite{KronigPenneydamped}. The tomograms of the quark states were introduces in \cite{ActaPhysHung}. 

The consideration of classical mechanics within the framework of the tomographic representation was done in \cite{OVMVIMLasRes1997}, \cite{2004JRLR}, \cite{JRLR2001}, \cite{hlimit}, \cite{Pilyavets}. The relation between the quantum state description and the classical state description is elucidated in \cite{Ibort}. The quantum--classical limits for quantum tomograms are studied and compared with the corresponding classical tomograms in \cite{Marmo3}. The inverted oscillator was discussed recently in \cite{Schleih,12a} in connection with relation to cosmological problems.  The inverted oscillator in the probability representation of quantum mechanics was considered in \cite{myEntropy2023}. Quantum evolution equation for open system was considered in   \cite{TeretenkovNosal,Teretenkov,ManciniPaulieq}. 
Schr\"odinger and von Neumann equations, as well as equations for the evolution of open systems, are written in the form of linear classical–like equations for the probability distributions determining the quantum system states in \cite{Entropy2021}. 
The evolution of tomograms for different quantum systems, both finite and infinite dimensional was considered in \cite{Kishore}. A method based on tomographic representation to simulate the quantum dynamics was applied to the wave packet tunneling of one and two interacting particles in \cite{Arhipov}.

The paper is organized as follows. In Sec.2 we provide definition of entangled probability distributions and describe some their properties. 
In Sec.3 we present the examples of entangled probability distributions of two random continious variables. The conclusions and prospectives are given in Sec.4.  

\section{Probability distributions}
Let us consider the set of continious random variables $\vec x=(x_1,x_2,\ldots, x_N)$ and parameter $\vec a =(a_1,a_2,\ldots, a_N)$ which can take some values and we consider a function of these  variables $0\leq w(\vec x|\vec a)\leq 1
$. This function satisfies the condition 
\[\int w(\vec x|\vec a)\prod_{k=1}^N d x_k=1.\]
We consider this nonnegative normalized function as a conditional probability distribution of $N$ random variables. Then we can introduce nonnegative functions $w_k(x_k|a_k)$ using the definition 
\begin{equation}\label{eq.1}
w_k(x_k|a_k)=\int w(\vec x|\vec a)\prod_{k'\neq k=1}^N d x'_k.
\end{equation}
We call the probability distribution $w(\vec x|\vec a)$ separable one if it equals to convex sum of the products of functions (\ref{eq.1})
\begin{equation}\label{eq.2}
w(\vec x|\vec a)=\sum_s{\cal P}_s\prod_{s=1}^N w_s(x_s|a_s),\quad {\cal P}_s\geq0,\quad \sum_s{\cal P}_s=1.
\end{equation}
If the function $w(\vec x|\vec a)$ can not be represented in the form of convex sum of the functions (\ref{eq.1}) we call the probability distribution $w(\vec x|\vec a)$ entangled probability distribution, i.e. 
\begin{equation}\label{eq.3}
w(\vec x|\vec a)\neq\sum_s{\cal P}_s\prod_{s=1}^N w_s(x_s|a_s).
\end{equation}
In the case of two variables $x_1,x_2$ and parameters $a_1,a_2$ we have for entangled probability distribution $w(x_1,x_2|a_1,a_2)$ the definition 
\begin{equation}\label{eq.4}
w(x_1,x_2|a_1,a_2)\neq\sum_s{\cal P}_s w_s(x_1|a_1)w_s(x_2|a_2).
\end{equation}

\section{Examples of the entangled probability distributions}
Let us remind the value of the propagator for the inverted oscillator evolution 
\begin{equation}\label{propagator} G(x,y,t)=\frac{1}{\sqrt{2\pi\sinh t}}\exp\left[ \frac{i}{2}\left(\coth t(y^2+x^2)-\frac{2x y}{\sinh t}\right)\right].
\end{equation}
We will use this formula later. In order to consider examples we will study properties of two mode oscillator state. We consider the very simple model of state $\psi(x,y)$ of the form  
\begin{equation}\label{eq.2.1}
\psi(x,y)=\frac{1}{\sqrt2}\left(\psi_0(x)\psi_1(y)+\psi_1(x)\psi_0(y)\right)=\frac{x+y}{\sqrt\pi}\exp\left(-\frac{x^2}{2}-\frac{y^2}{2}\right).\end{equation}
The function (\ref{eq.2.1}) is the superposition of the wave functions of one-mode oscillators, where first function is ground state of the first oscillator
\begin{equation}\label{eq.2.2}
\psi_0(x)=\frac{e^{-\frac{x^2}{2}}}{\pi^{1/4}}
\end{equation}
and the second function is the first excited state of the second oscillator
\begin{equation}\label{eq.2.3}
\psi_1(y)=\frac{\sqrt2y}{\pi^{1/4}}e^{-\frac{y^2}{2}}.
\end{equation}
The normalization constant is such that we have
\begin{equation}\label{eq.2.4}
\int_{-\infty}^\infty\psi(x,y)\psi^\ast(x,y) d x d y=1.
\end{equation}
We calculated this normalization condition using the known formula 
\begin{equation}\label{eq.2.5}
\int_{-\infty}^\infty\exp\left(-a x^2+b x\right)d x=\sqrt{\frac{\pi}{a}}\exp\left(\frac{b^2}{4a}\right).
\end{equation}
Considering the derivative of both sides of the above equality with respect to the parameter $b$ we obtain the integral relation 
\begin{equation}\label{eq.2.6}
\int_{-\infty}^\infty x\exp\left(-a x^2+b x\right)d x=\sqrt{\frac{\pi}{a}}\frac{b}{2 a}\exp\left(\frac{b^2}{4a}\right).
\end{equation}
To calculate another integral we calculate derivative of the both sides of equality (\ref{eq.2.5}) with respect to the parameter $-a$ then we get the value of the integral   
\begin{equation}\label{eq.2.7}
\int_{-\infty}^\infty x^2\exp\left(-a x^2+b x\right)d x=\frac{1}{2a}\sqrt{\frac{\pi}{a}}\exp\left(\frac{b^2}{4a}\right)\left(1+\frac{b^2}{2a}\right).\end{equation}
Using the values of calculating integrals we get the normalization condition
\begin{equation}\label{eq.2.8}
\frac{1}{\pi}\int_{-\infty}^\infty(x+y)^2\exp\left(-x^2-y^2\right)d x d y=1.
\end{equation}
To consider the map of the wave function $\psi(x,y)$ onto the probability distribution let us study the integral
\begin{equation}\label{eq.2.10}
A=\int_{-\infty}^\infty\psi(x,y)\exp\left[i\left(\frac{\mu_1x^2}{2\nu_1}-\frac{X x}{\nu_1}+\frac{\mu_2y^2}{2\nu_2}-\frac{Y y}{\nu_2}\right)\right]d x d y.
\end{equation}
The integrals (\ref{eq.2.5})-(\ref{eq.2.7}) give the possibility to obtain the relation 
\begin{eqnarray}
&&A=\frac{\sqrt\pi}{\sqrt{\left(1-\frac{i\mu_1}{\nu_1}\right)\left(1-\frac{i\mu_2}{\nu_2}\right)}}
\left(-\frac{2i X}{\nu_1\left(1-\frac{i\mu_1}{\nu_1}\right)}-\frac{2i Y}{\nu_2\left(1-\frac{i\mu_2}{\nu_2}\right)}\right)\nonumber\\
&&\times\exp\left[-\frac{X^2}{2\nu_1^2\left(1-\frac{i\mu_1}{\nu_1}\right)}
-\frac{Y^2}{2\nu_2^2\left(1-\frac{i\mu_2}{\nu_2}\right)}\right].\label {eq.2.11}
\end{eqnarray}
The probability distribution related to these formulae reads (see, for example, \cite{Entropy2021}) 
\begin{equation}\label{eq.2.12}
w=\frac{A A^\ast}{4\pi^2|\nu_1||\nu_2|}.
\end{equation}
In explicit form the probability distribution $w(X,Y|\mu_1,\nu_1,\mu_2,\nu_2)$ is given by the relation 
\begin{eqnarray}
w(X,Y|\mu_1,\nu_1,\mu_2,\nu_2)=&&\frac{\left(\nu_2^2+\mu_2^2\right)X^2+2\left(\nu_1\nu_2+\mu_1\mu_2\right)X Y+\left(\nu_1^2+\mu_1^2\right) Y^2}{\pi \left(\nu_1^2+\mu_1^2\right)^{3/2}\left(\nu_2^2+\mu_2^2\right)^{3/2}}\nonumber\\
&&\times\exp\left[-\frac{X^2}{\mu_1^2+\nu_1^2}-\frac{Y^2}{\mu_2^2+\nu_2^2}\right].\label{eq.2.13}
\end{eqnarray}
For particular case $\nu_1=\nu_2=1,\quad \mu_1=\mu_2=0,$ one gets 
\begin{equation}\label{eq.2.14} w(X,Y|\mu_1=0,\nu_1=1,\mu_2=0,\nu_2=1)=\frac{1}{\pi}(X+Y)^2\exp\left(-X^2-Y^2\right). 
\end{equation}
One can check, that the function $w(X,Y,|\mu_1,\nu_1,\mu_2,\nu_2)$ (\ref{eq.2.14}) satisfy the condition 
\begin{equation}\label{eq.2.15}
\int\int w(X,Y,|\mu_1,\nu_1,\mu_2,\nu_2) d X d Y=1.
\end{equation}
As we know, this probability distribution function determines the quantum states which are entangled states. Due to this we call this probability distribution entangled probability distribution. In quantum mechanics the wave functions of two--mode oscillators which are obtained by means of superposition of two different wave functions are entangled pure states. In connection with this the tomographic probability distribution is described by the probability distribution function (\ref{eq.2.13}) and it cannot be represented in the form of equation (\ref{eq.2}). On the other hand the integral 
\begin{equation}\label{eq.2.16}
w(X|\mu_1,\nu_1)=\int w(X,Y|\mu_1,\nu_1,\mu_2,\nu_2)d Y=\frac{\exp\left(-\frac{X^2}{\mu_1^2+\nu_1^2}\right)}{\sqrt{\pi\left(\mu_1^2+\nu_1^2\right)}}
\left[\frac{1}{2}+\frac{X^2}{\mu_1^2+\nu_1^2}
\right]. 
\end{equation}
One can check, that 
\begin{equation}\label{eq.2.17}
\int w(X|\mu_1,\nu_1)d X=1.
\end{equation}
This function $w(X|\mu_1,\nu_1)$ is marginal conditional probability distribution of position $X$ which is the position of the first oscillator and the conditions are labeled by the real parameters $\mu_1$, $\nu_1$. Also, if we repeat analogous calculations for the second oscillator we will get  
\begin{equation}\label{eq.2.18}
w(Y|\mu_2,\nu_2)=\int w(X,Y|\mu_1,\nu_1,\mu_2,\nu_2)d X=\frac{\exp\left(-\frac{Y^2}{\mu_2^2+\nu_2^2}\right)}{\sqrt{\pi\left(\mu_2^2+\nu_2^2\right)}}
\left[\frac{1}{2}+\frac{Y^2}{\mu_2^2+\nu_2^2}.
\right]. 
\end{equation}
One can check, that 
\begin{equation}\label{eq.2.19}
\int w(Y|\mu_2,\nu_2)d Y=1.
\end{equation}
This function $w(Y|\mu_2,\nu_2)$ is marginal conditional probability distribution of position $Y$ which is the position of the second  oscillator and the conditions are labeled by the real parameters $\mu_2$, $\nu_2$.

Let us calculate other probability distributions given by two density operators ${\hat{\bar\rho}}_0$, $
{\hat{\widetilde 
\rho}}_1$ 
\begin{equation}\label{eq.2.20}
\hat\rho(0,1)=\frac{1}{2}\left({\hat{\bar\rho}}_0 {\hat{\widetilde 
		\rho}}_1+{\hat{\bar\rho}}_1 {\hat{\widetilde 
		\rho}}_0\right), 
\end{equation}
here ${\hat{\bar\rho}}_0(x,x')=\psi_0( x)\psi_0^\ast(x')$ and 
${\hat{\widetilde\rho}}_1(y,y')=\psi_1(y)\psi^\ast_1(y') $. This   density operator $\hat \rho(0,1)$ describes the separable states of the two-mode oscillator. Other two separable states are described by operators ${\hat{\bar\rho}}_0 {\hat{\widetilde 
	\rho}}_1$ and ${\hat{\bar\rho}}_1 {\hat{\widetilde 
\rho}}_0$. The state with probability distribution (\ref{eq.2.13}) 
is the entangled state. The three states which are separable states are described by the probability distributions 
\begin{eqnarray}\label{last1}
w_{\hat \rho(0,1)}(X,Y|\mu_1,\nu_1,\mu_2,\nu_2)
=\frac{\exp\left(-\frac{X^2}{\mu_1^2+\nu_1^2}-\frac{Y^2}{\mu_2^2+\nu_2^2}\right)}{\pi\left(\mu_1^2+\nu_1^2\right)^{1/2}\left(\mu_2^2+\nu_2^2\right)^{1/2}}\left(\frac{Y^2}{\mu_2^2+\nu_2^2}+\frac{X^2}{\mu_1^2+\nu_1^2}\right),
\end{eqnarray} 
\begin{equation}\label{last2}
w_{{\hat{\bar\rho}}_0 {\hat{\widetilde\rho}}_1}(X,Y|\mu_1,\nu_1,\mu_2,\nu_2)=
\frac{2Y^2}{\pi \left(\mu_1^2+\nu_1^2\right)^{1/2}\left(\mu_2^2+\nu_2^2\right)^{3/2}}
\exp \left(-\frac{X^2}{\mu_1^2+\nu_1^2}-\frac{Y^2}{\mu_2^2+\nu_2^2}\right),
\end{equation}
\begin{equation}\label{last3}
w_{{\hat{\bar\rho}}_1 {\hat{\widetilde 
		\rho}}_0}(X,Y|\mu_1,\nu_1,\mu_2,\nu_2)=\frac{2X^2}{\pi \left(\mu_2^2+\nu_2^2\right)^{1/2}\left(\mu_1^2+\nu_1^2\right)^{3/2}}
	\exp \left(-\frac{X^2}{\mu_1^2+\nu_1^2}-\frac{Y^2}{\mu_2^2+\nu_2^2}\right).
\end{equation}
Comparing (\ref{eq.2.13}) with the (\ref{last1}), (\ref{last2}), {\ref{last3}) we see that the entangled state probability distribution (\ref{eq.2.13})  is essentially different from the separable state probability distributions (\ref{last1}), (\ref{last2}), {\ref{last3}). 
		
		Let us calculate the marginal distributions for the probability distributions given by the formulae 
	(\ref{last1}), (\ref{last2}), ({\ref{last3}). We get  
		\begin{eqnarray}
		&&w_{\hat \rho(0,1)}(X|\mu_1,\nu_1)=\int w_{\hat \rho(0,1)}(X,Y|\mu_1,\nu_1,\mu_2,\nu_2) d Y=\nonumber\\
		&&\frac{1}{2\sqrt{\pi\left(\nu_1^2+\mu_1^2\right)}}
		\exp\left(-\frac{X^2}{\mu_1^2+\nu_1^2}\right)\left(1+\frac{2X^2}{\mu_1^2+\nu_1^2}\right), \label{last4} 
		\end{eqnarray}	
			\begin{eqnarray}
		&&w_{\hat \rho(0,1)}(Y|\mu_2,\nu_2)=\int w_{\hat \rho(0,1)}(X,Y|\mu_1,\nu_1,\mu_2,\nu_2) d X=\nonumber\\
		&&\frac{1}{2\sqrt{\pi\left(\nu_2^2+\mu_2^2\right)}}
		\exp\left(-\frac{Y^2}{\mu_2^2+\nu_2^2}\right)\left(1+\frac{2Y^2}{\mu_2^2+\nu_2^2}\right), \label{last5} 
		\end{eqnarray}	
			\begin{eqnarray}
		&&
		w_{{\hat{\bar\rho}}_0 {\hat{\widetilde\rho}}_1}(X|\mu_1,\nu_1)=\int w_{{\hat{\bar\rho}}_0 {\hat{\widetilde\rho}}_1}(X,Y|\mu_1,\nu_1,\mu_2,\nu_2) d Y=\nonumber\\
	&&
		\frac{1}{\sqrt{\pi\left(\nu_1^2+\mu_1^2\right)}}
		\exp\left(-\frac{X^2}{\mu_1^2+\nu_1^2}\right), \label{last6} 
		\end{eqnarray}	
		
			\begin{eqnarray}
		&&
		w_{{\hat{\bar\rho}}_0 {\hat{\widetilde\rho}}_1}(Y|\mu_2,\nu_2)=\int w_{{\hat{\bar\rho}}_0 {\hat{\widetilde\rho}}_1}(X,Y|\mu_1,\nu_1,\mu_2,\nu_2) d X=\nonumber\\
		&&
	\frac{2Y^2}{\sqrt{\pi}\left(\nu_2^2+\mu_2^2\right)^{3/2}}
		\exp\left(-\frac{Y^2}{\mu_2^2+\nu_2^2}\right),
		\label{last7} 
		\end{eqnarray}	
		
			\begin{eqnarray}
	&&
	w_{{\hat{\bar\rho}}_1 {\hat{\widetilde 
				\rho}}_0}(Y|\mu_2,\nu_2)=\int w_{{\hat{\bar\rho}}_1 {\hat{\widetilde 
				\rho}}_0}(X,Y|\mu_1,\nu_1,\mu_2,\nu_2) d X=\nonumber\\
	&&
	\frac{1}{\sqrt{\pi\left(\nu_2^2+\mu_2^2\right)}}
	\exp\left(-\frac{Y^2}{\mu_2^2+\nu_2^2}\right), \label{last8} 
	\end{eqnarray}	
	
	\begin{eqnarray}
	&&
	w_{{\hat{\bar\rho}}_1 {\hat{\widetilde 
				\rho}}_0}
	(X|\mu_1,\nu_1)=\int w_{{\hat{\bar\rho}}_1 {\hat{\widetilde 
				\rho}}_0}(X,Y|\mu_1,\nu_1,\mu_2,\nu_2) d Y=\nonumber\\
	&&
	\frac{2X^2}{\sqrt{\pi}\left(\nu_1^2+\mu_1^2\right)^{3/2}}
	\exp\left(-\frac{X^2}{\mu_1^2+\nu_1^2}\right),
	\label{last9} 
	\end{eqnarray}		
Comparing (\ref{eq.2.16}) and (\ref{eq.2.18}) with  (\ref{last4})-(\ref{last9}) we also see the difference of marginal distributions in entangled state (\ref{eq.2.13}) and in separable states (\ref{last1}), (\ref{last2}), (\ref{last3}). There exists a central mass tomography of two--mode oscillator \cite{center,center1,center2}.  These oscillator states are described by the probability distributions of one random variable (the center of mass position). We will study entangled and separable states of the oscillator in future publications.

\section{Conclusion}	
We conclude mentioning the main ideas to be used on the bases of the probability representation of quantum mechanics. In fact this approach provides the possibility to formulate new aspects of probability theory and group representation theory which play important role in quantum mechanics foundations. As we know the Jordan–Schwinger map \cite{Jordan,schwinger} provides the possibility to map the properties of matrices like the $N\times N$ matrices $A$, $B$, $C$ with commutation relation $[A,B]=C$ onto three operators $\hat A$, $\hat B$, $\hat C$ acting in the Hilbert space of $N$-dimensional oscillator states, namely, $A\leftrightarrow \hat A=\sum_{i, k=1}^N A_{i k}\hat a^\dagger_i\hat a_k$,  $B\leftrightarrow \hat B=\sum_{i, k=1}^N B_{i k}\hat a^\dagger_i\hat a_k$, $C\leftrightarrow \hat C=\sum_{i, k=1}^N C_{i k}\hat a^\dagger_i\hat a_k$, where $[\hat a_i,\hat a^\dagger_k]=\delta_{i k}$ and $\hat a_i$, $\hat a^\dagger_k$ are the annihilation and creation boson operators of the $N$-mode oscillator. The commutator $[\hat A,\hat B]=\hat C$ of the oscillator operators repeats the commutator properties of the matrices $A$, $B$, $C$. For example, the spin Pauli matrices commutation relations are reproduced by the  two-dimensional oscillator bosonic commutation relations and the states of $s$-spin are mapped onto the excited energy states of the energy levels of the oscillator. We get the results that spin states can be invertibly mapped onto energy levels states of the two-mode oscillator. From this observation follows an idea that all the spin states can be mapped on the probability distributions describing the states of the two-mode oscillator (or the Hermitte polynomials  \cite{yesterday1,yesterday2,yesterday3}). This provides the possibility to map all the spin states (irreducible representations  of rotation $SU(2)$ group) onto transforms of the probability distributions which is new aspect of the group representation theory. If to apply this idea to $N$-dimensional oscillator properties we extend the result to other group representation theory ($SU(N)$ group and for other Lie groups). The $s$-spin states can be described as the multi--mode states associated with probability distributions of multidimensional Hermite  polynomials. The approach will be considered in future publications.

\end{document}